OpenStudy (anonymous):

If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4). 17 −17 5 −5

OpenStudy (anonymous):

(f-g)(4) = f(4) - g(4) so find f(4) and g(4) then subtract....

OpenStudy (anonymous):

stilll need help

jhonyy9 (jhonyy9):

so than f(4)=2*4-2 and g(4)= this will make you ok ?

OpenStudy (anonymous):

ok so iz iht -5

hero (hero):

(f-g)(4) = f(4) - g(4) = 2(4) - 2 - (3(4) - 1) = 8 - 2 - 12 + 1 = 8 + 1 - 2 - 12 = 9 - 14 = -5

OpenStudy (anonymous):

thanx